How do you solve a square and a rectangle with the same perimeter? They give us 4x+3 on the square for one side and 5x on the rectangle/
Answer by stanbon(75887)

How do you solve a square and a rectangle with the same perimeter? They give us 4x+3 on the square for one side and 5x on the rectangle/
-------
If the side of the square is 4x+3, the perimeter is 4(4x+3) = 16x + 12
-------
If one side of the rectangle is 5x and the perimeter is 16x+12,
p = 2(L + W)
16x+12 = 2(5x + W)
8x + 6 = 5x + W
w = 3x+6
---------
Dimensions of the rectangle are 5x and 3x+6.
====================
Chees,
Stan H.
====================